# The Quantum Measurement Problem: A Review of Recent Trends

### v2

@@ Abstract
Keywords Decoherence; Measurement Problem; Quantum state collapse; Quantum mechanics interpretation
@@ 1 Introduction
Quantum mechanics, as we learn in college, posits that an isolated system evolves unitarily and deterministically following the Schrödinger Equation. When we measure the system, the Born rule determines the output probabilities. The quantum state collapses in a non-unitary, stochastic way, so the same outcome is obtained if the measurement is repeated. Despite its awkward double-evolution standard, this seemingly simple set of rules enabled comprehension of nature and technological advances to a level our species has never seen.
@@
Nevertheless, let us ask what “measure the system” means. We will not find an answer in the standard quantum mechanics formulation. This innocent question unravels a disturbing thread of conceptual problems, which are surprisingly difficult to address experimentally and may have dramatic philosophical consequences for our conception of the world.
@@
The Measurement Problem has been haunting physicists and philosophers of science for almost as long as the century since quantum mechanics was proposed. We can take as an anecdotal example the World Science Festival held in the first semester of 2024, a science popularization event hosted by Brian Greene. Greene asked three quantum mechanics specialists about the Quantum Measurement Problem and got three entirely distinct answers. The entry Interpretations of Quantum Mechanics in Wikipedia contains thirteen distinct classes of theories and interpretations, and it does not even account for some recent proposals.
@@
The cacophony of proposed solutions to the Measurement Problem—including voices from parts of the physics community that question whether the problem even exists—is disturbing. It hurts our conception of science as delivering a consistent description of reality. Why, after a century of quantum mechanics, do we still have no consensus on its most fundamental process? What do we truly understand about the Quantum Measurement Problem?
@@
Any newcomer trying to get an answer to these questions faces a formidable challenge: the Measurement Problem research is spread throughout countless isolated sub-communities. These groups don’t dialogue with each other, making it exceptionally difficult to build an overview of what is going on. The large number of single-author papers doesn’t help either. These papers raise hypotheses and propose theories and tests. However, most don’t resonate within a mature community and only add to the cacophony.
@@
As newcomers in the field ourselves—we just published our first contribution recently [1]—we felt firsthand all these issues. We tackled this challenge with extensive reading, discussion, and consultation with specialists on specific topics. We explored all corners of the field rather than focusing on a single subfield. This review arose as a natural consequence. We aim to deliver a broad overview of the several competing research lines. We don’t aim at a comprehensive survey (which should not be possible within reasonable limits of a paper) but to illuminate the main questions being currently discussed, allowing curious non-specialists to start digging the literature by themselves. In doing so, we hope to show that the Measurement Problem is not as bleak as it seems, although there is still much to uncover.
@@
Given the Measurement Problem’s central role, it’s unsurprising that several other reviews have been recently published to explore its unresolved aspects from different perspectives. Hance and Hossenfelder [2] critically assess whether the problem is fundamental or an artifact of our theoretical framework. In a synthetic review, they identify five requirements that any satisfactory solution to the problem must satisfy and analyze several proposed solutions. Müller [3] followed a different path. Instead of focusing on the proposed solutions, he approaches the measurement problem from a formal logic perspective. He decomposes it into six independent problems, which should be investigated separately. Meanwhile, Freire Jr. [4] delivers a historical account, linking the foundations of quantum mechanics to current scientific and technological advances.
@@
Our work also brings a distinct perspective by examining the Measurement Problem through the chemistry lens (since two authors, RSM and MB, are theoretical chemists). Historically, chemists have treated quantum mechanics pragmatically, using it as a tool rather than engaging with its foundational questions. However, recent advances in quantum materials [10], quantum dots for imaging and drug delivery [11], ultrafast spectroscopy [12], and quantum computing [13, 14] have brought the Quantum Measurement Problem into sharper focus at the chemistry labs. These fields increasingly rely on quantum coherence, entanglement, and decoherence—concepts central to quantum measurement theory [15]. As chemists push the boundaries of quantum state control, they may need to reconsider long-standing assumptions about wave function collapse, observer effects, and the role of decoherence in chemical processes.
Despite being unconventional in scientific papers, we extensively referred to non-peer-reviewed web resources, including books, essays, and videos. These resources offer a plain language introduction and analysis without intimidating technicalities. They naturally don’t replace technical literature, but should not be neglected as relevant entry points to the subject.
@@ 2 Quantum State Description and Evolution
Consider, for example, the ammonia molecule NH3 (Figure 1). Its more stable geometry resembles a pyramid, with the three hydrogen atoms forming a triangular basis and the nitrogen atom lying above or below it. Thus, ignoring molecular vibrations and rotations, the nuclear quantum state of ammonia has two components, corresponding to pyramids up and down [20, Chs 8 & 9]. Let’s name them and states, respectively, meaning we will treat NH3 as a quantum bit (or a qubit).
@@
where each and is a density operators on and , and , . Otherwise, is entangled.
@@
In terms of the density operator, this probability becomes
@@
According to the standard postulates of quantum mechanics, after the measurement, the system is left in the normalized eigenstate if was measured. This state change is referred to as quantum state collapse. In this discussion, we assumed that collapse is instantaneous and occurs at . Standard quantum theory gives no clue on whether this is precise or not. Recent resonant inelastic X-ray scattering (RIXS) experiments have suggested that state collapse occurs gradually over a finite time [21]. However, this interpretation remains ambiguous, as the observed dynamics could also be explained by progressive decoherence rather than a fundamental modification of quantum measurement. We will return to when and how long it takes for collapse to occur in discussing Objective Collapse Theories in Section 5.3. The implications of an instantaneous collapse are explored in Section 5.7, in the context of quantum field measurements.
@@ 3 The Problems with Quantum Measurements
To summarize the description of the quantum state evolution discussed in the previous section, standard quantum mechanics prescribes that a quantum state follows two types of time evolution: Left on itself, it evolves with the Schrödinger Equation, which is unitary and deterministic [16, Ch 3][17, Ch 1]. During a measurement, the system produces a classical outcome with probabilities given by the Born rule. In idealized cases, this is often described by a non-unitary state update (the projection postulate), but more generally, a measurement is any transformation that extracts classical information from a quantum system. These processes must be probabilistic and cannot allow communication between distant systems without a physical carrier.
There are different ways to scrutinize this description [3, 22, 23]. Here, we follow Schlosshauer [24, Ch 2], who decomposed it into three problems: the problem of the preferred basis, the problem of the nonobservability of interference, and the problem of outcomes. Each of these challenges arises naturally from the standard formulation of quantum mechanics and highlights a gap in our understanding of what happens when a quantum system is measured. Below, we define these problems in sequence.
@@
This ambiguity, why a specific set of outcomes is realized when many are in principle available, is the problem of the preferred basis, first formulated by Zurek [25].
The problem of the nonobservability of interference. Quantum mechanics predicts that systems can exist in coherent superpositions, which should produce observable interference effects. Indeed, such effects are routinely seen at microscopic scales—for instance, in electron double-slit experiments [26] or the oscillations of generalized oscillator strengths for symmetric molecules under electron impact [27]. More recently, Kanitz and co-workers demonstrated interference patterns from helium and hydrogen atoms transmitted through single-layer graphene, achieving diffraction with atoms at kiloelectronvolt energies [28].
@@ 4 Decoherence
Because of the system-environment interaction, the quantum state evolves into an entanglement:
@@
In general, the environmental states are not orthogonal, and the coherences are non-null. However, as time passes, the environment states tend to become orthogonal because their dynamics are different when the environment interacts with the system’s and geometries. Thus, the off-diagonal terms tend to zero,
@@
In formal terms, decoherence arises from the time propagation of the reduced density matrix, which evolves with a master equation of the type [24, Ch 4].
where is the system’s Hamiltonian, including the environment’s perturbation, and it is the term responsible for decoherence to the environment.
@@
Although decoherence typically occurs in femtoseconds or even sub-femtoseconds, specific molecular systems exhibit surprisingly long-lived coherence effects. For instance, Kaufman et al. [32] demonstrated that electronic coherences could survive for hundreds of femtoseconds in molecules with parallel potential energy surfaces due to reduced dephasing between branched wavepackets (see also 4). Schultz et al. [15] comprehensively review cases where vibronic coupling and structured environments lead to long coherence lifetimes, including examples from light-harvesting complexes in photosynthesis.
Babcock et al. [33] reported superradiant effects in tryptophan networks extending over microtubule-scale architectures, which require long-lived electronic coherence. This is remarkable, since superradiance typically unfolds over hundreds of femtoseconds [34]. For isolated tryptophan in aqueous solution, electronic decoherence occurs in less than 100 fs [35], implying that the protein scaffold must be critical in preserving coherence to allow superradiant emission. Babcock’s result has also attracted media attention (see here and here) because it may support Hammeroff-Penrose’s hypothesis on the quantum origin of consciousness (see Section 5.6.2).
@@
In addition to suppressing quantum interferences, decoherence also addresses the problem of the preferred basis. It leads to the selection of a stable outcome basis through a process known as einselection, which drives the quantum state to the basis least entangled with the environment (this basis is known as the pointer basis) [25]. This preferred basis depends on the Hamiltonian [37]. Assuming that the position operator commutes with the system-environment interaction Hamiltonian (usually the case), decoherence tends to select the position eigenstates as the preferred basis as long as the interaction Hamiltonian is more important for the dynamics than the system’s Hamiltonian. If both terms compete, the preferred basis is close to coherent states, which minimizes the position and momentum uncertainty. If the system’s Hamiltonian dominates the dynamics, energy eigenstates are preferred as the basis.
For a molecule in a gas, Paz and Zurek’s analysis [37] implies that the outcome basis of the environment selects for measuring nuclei is approximately position states. On the other hand, energy states emerge as the preferred outcome basis for measuring electrons. Thus, when we initially chose the basis for describing the quantum state of the ammonia molecule, we did it because it matched our classical intuition of nuclei corresponding to particles with defined positions. After discussing decoherence, we realize that we may have developed such a classic intuition because the environment selects the basis and suppresses superposition effects, creating the world of particles with well-defined positions.
Nevertheless, einselection may not be the last word in the preferred basis problem. Adil and co-authors have recently shown that the preferred basis is not solely dictated by the system-environment interaction but also by the natural factorization of the global Hilbert space [38]. Different factorizations can lead to multiple coexisting sets of pointer states, meaning classical behavior emerges in different ways depending on the structure of the Hamiltonian. These coexisting classical realms challenge the idea that einselection always yields a single classical outcome. They may even reignite the preferred basis problem, but now at the level of Hamiltonian factorization.
@@
This dominant role of decoherence in the current understanding of quantum mechanical processes doesn’t imply it is a consensus. Landsmann, for instance, characterizes decoherence as “an unmitigated disaster” [39, Ch 11]. His core criticism is that decoherence rigorously only occurs in an environment with infinite degrees of freedom after infinite interaction time with the system. Kastner reinforces this criticism by noticing that decoherence claims of describing irreversible processes suffer from similar conceptual problems already present in statistical mechanics when explaining the emergence of irreversibility from reversible processes [40].
@@ 5 Facing the Problem of Outcomes
In the previous section, we have seen that decoherence clarifies a significant portion of the Measurement Problem. However, the problem of outcomes is still open: After the basis is chosen and quantum superpositions are suppressed, the system still remains in a mixture of possible outcomes. Decoherence does not tell how and why only one of these outcomes is measured. At this point, standard quantum mechanics interpretation claims the occurrence of an additional dynamic step, the collapse:
We illustrate these different processes with an example of a photoexcited molecule undergoing internal conversion in Figure 4. Light initially promotes the molecule to an excited electronic state. The nuclear wavepacket evolves on the potential energy surface of the excited state until it reaches a degeneracy region with the ground state potential energy surface. There, quantum coherence builds between the two states, creating a state superposition. The dephasing between the wavepacket components in each state counteracts, causing decoherence. All events up to this point, including decoherence, can be simulated by propagating the Schrödinger equation for the molecule [41]. If the molecule is measured, it will either be in the ground or excited state. However, our simulation can only tell the probabilities of these outputs.
@@
In the following subsections, we will dive into a few classes of potential solutions that represent the current debate on the Measurement Problem, with little attention to philosophical and historical aspects. More comprehensive approaches can be found in Refs.[4], [39, Introduction & Ch 11], [43, 44].
@@
As stated in Eq. (25), decoherence acts on the off-diagonal terms of the system’s reduced density matrix. Collapse, however, zeroes the diagonal terms except for one element of the entire density matrix (including the observer), which is not represented in Eq. (25). We can be a bit more flexible with this definition. In principle, a quantum theory including collapse could go as far as to provide a diagonal density matrix. This means we would be left in a situation of statistical uncertainty: the collapse occurred, but the theory only informed us of the probability of each outcome. We could call such a case an epistemic collapse. If, however, the theory zeroes the diagonal (except one) elements, then we would have an ontic collapse.

### v3

@@ Abstract
Keywords: decoherence; measurement problem; quantum state collapse; interpretations of quantum mechanics
Contents
- 1 Introduction
- 2 Quantum State Description and Evolution
- 3 The Problems with Quantum Measurements
- 4 Decoherence
- 5 Facing the Problem of Outcomes
- 6 The Ontological Status of the Wave Function
- 7 Through the Forest of Quantum Foundations
@@ 1 Introduction
Quantum mechanics, as we learn in college, posits that an isolated system evolves unitarily and deterministically following the Schrödinger equation. When we measure the system, the Born rule determines the output probabilities. The quantum state collapses in a non-unitary, stochastic way, so the same outcome is obtained if the measurement is repeated. Despite its awkward double rule of evolution, this seemingly simple framework enabled comprehension of nature and technological advances to a level that our species has never seen before.
@@
Nevertheless, let us ask what “measure the system” means. We will not find an answer in the standard formulation of quantum mechanics. This innocent question unravels a disturbing thread of conceptual problems, which are surprisingly difficult to address experimentally and may have dramatic philosophical consequences for our conception of the world.
@@
The Measurement Problem has been haunting physicists and philosophers of science for almost as long as the century since quantum mechanics was proposed. We can take as an anecdotal example the World Science Festival held in the first semester of 2024, a science popularization event hosted by Brian Greene. Greene asked three quantum mechanics specialists about the Measurement Problem and got three entirely distinct answers. The entry Interpretations of Quantum Mechanics in Wikipedia contains thirteen distinct classes of theories and interpretations, and it does not even account for some recent proposals.
@@
The cacophony of proposed solutions to the Measurement Problem—including voices from parts of the physics community that question whether the problem even exists—is disturbing. It hurts our conception of science as delivering a consistent description of reality. Why, after a century of quantum mechanics, do we still have no consensus on its most fundamental process? What do we truly understand about the Measurement Problem?
@@
Any newcomer trying to get an answer to these questions faces a formidable challenge: the Measurement Problem research is spread throughout countless isolated sub-communities. These groups do not talk to each other, making it exceptionally difficult to build an overview of what’s going on. The large number of single-author papers doesn’t help either. These papers raise hypotheses and propose theories and tests. However, most don’t resonate within a mature community and only add to the cacophony.
@@
As newcomers to the field ourselves—we just published our first contribution recently [1]—we felt firsthand all these issues. We tackled this challenge with extensive reading, discussion, and consultation with specialists on specific topics. We explored all corners of the field rather than focusing on a single subfield. This review arose naturally. We aim to deliver a broad overview of the several competing research lines. We don’t aim for a comprehensive survey (which would not be possible within the reasonable limits of a paper) but rather to illuminate the central questions currently being discussed, allowing curious non-specialists to start exploring the literature on their own. In doing so, we hope to show that the Measurement Problem is not as bleak as it seems, although there is still much to be uncovered.
@@
Given the Measurement Problem’s central role, it is unsurprising that several other reviews have been recently published to explore its unresolved aspects from different perspectives. Hance and Hossenfelder [2] critically assess whether the problem is fundamental or an artifact of our theoretical framework. In a synthetic review, they identify five requirements that any satisfactory solution to the problem must satisfy and analyze several proposed solutions. Müller [3] followed a different path. Instead of focusing on the proposed solutions, he approaches the Measurement Problem from a formal logic perspective. He decomposes it into six independent problems, which should be investigated separately. Meanwhile, Freire Jr. [4] delivers a historical account, linking the foundations of quantum mechanics to current scientific and technological advances.
@@
Our work also brings a distinct perspective by examining the Measurement Problem through the chemistry lens (since two authors, RSM and MB, are theoretical chemists). Historically, chemists have treated quantum mechanics pragmatically, using it as a tool rather than engaging with its foundational questions. However, recent advances in quantum materials [10, 11], quantum dots for imaging and drug delivery [12], ultrafast spectroscopy [13], and quantum computing [14, 15] have brought the Measurement Problem into sharper focus at the chemistry labs. These fields increasingly rely on quantum coherence, entanglement, and decoherence—concepts central to quantum measurement theory [16, 17]. As chemists push the boundaries of quantum state control, they may need to reconsider long-standing assumptions about wave function collapse, observer effects, and the role of decoherence in chemical processes.
Despite being unconventional in scientific papers, we extensively referred to non-peer-reviewed web resources, including books, essays, and videos. These resources offer a plain language introduction and analysis free from intimidating technicalities. They naturally don’t replace technical literature, but should not be neglected as relevant entry points to the subject.
@@ 2 Quantum State Description and Evolution
Consider, for example, the ammonia molecule NH3 (Figure 1). Its more stable geometry resembles a pyramid, with the three hydrogen atoms forming a triangular basis and the nitrogen atom lying above or below it. Thus, ignoring molecular vibrations and rotations, the nuclear quantum state of ammonia has two components, corresponding to pyramids up and down [22, Chs 8 & 9]. Let’s name them and states, respectively, meaning we will treat NH3 as a quantum bit (or a qubit).
@@
where each and are density operators on and , and , . Otherwise, is entangled.
@@
where is the corresponding eigenstate. In terms of the density operator, this probability becomes
@@
According to the standard postulates of quantum mechanics, after the measurement, the system is left in the normalized eigenstate if was measured. This state change is referred to as quantum state collapse. In this discussion, we assumed that collapse is instantaneous and occurs at . Standard quantum theory gives no clue as to whether this is precise or not. Recent resonant inelastic X-ray scattering (RIXS) experiments have suggested that state collapse occurs gradually over a finite time [23]. However, this interpretation remains ambiguous, as the observed dynamics could also be explained by progressive decoherence rather than a fundamental modification of quantum measurement. We will return to when and how long it takes for collapse to occur in discussing Objective Collapse Theories in Section 5.3. The implications of an instantaneous collapse are explored in Section 5.7, in the context of quantum field measurements.
@@ 3 The Problems with Quantum Measurements
To summarize the description of the quantum state evolution discussed in the previous section, standard quantum mechanics prescribes that a quantum state follows two types of time evolution: Left on its own, it evolves with the Schrödinger equation, which is unitary and deterministic [18, Ch 3][19, Ch 1]. During a measurement, the system produces a classical outcome with probabilities given by the Born rule. In idealized cases, this is often described by a non-unitary state update (the projection postulate), but more generally, a measurement is any transformation that extracts classical information from a quantum system. These processes must be probabilistic and cannot allow communication between distant systems without a physical carrier.
There are different ways to scrutinize this description [3, 24, 25]. Here, we follow Schlosshauer [26, Ch 2], who decomposed it into three problems: the problem of the preferred basis, the problem of the nonobservability of interference, and the problem of outcomes. Each of these challenges arises naturally from the standard formulation of quantum mechanics and highlights a gap in our understanding of what happens when a quantum system is measured. Below, we define these problems in sequence.
@@
This ambiguity—why a specific set of outcomes is realized when many are in principle available—is the problem of the preferred basis, first formulated by Zurek [27].
The problem of the nonobservability of interference. Quantum mechanics predicts that systems can exist in coherent superpositions, which should produce observable interference effects. Indeed, such effects are routinely seen at microscopic scales—for instance, in electron double-slit experiments [28] or the oscillations of generalized oscillator strengths for symmetric molecules under electron impact [29]. More recently, Kanitz and coworkers demonstrated interference patterns from helium and hydrogen atoms transmitted through single-layer graphene, achieving diffraction with atoms at kiloelectronvolt energies [30].
@@ 4 Decoherence
Because of the interaction between the system and its environment, the joint quantum state becomes entangled:
@@
In general, the environmental states are not orthogonal, and the coherences are non-null. However, as time passes, the environment states tend to become effectively orthogonal because their dynamics are different when the environment interacts with the system’s and geometries. Thus, the off-diagonal terms tend to zero,
@@
In formal terms, decoherence arises from the time propagation of the reduced density matrix, which evolves with a master equation of the type [26, Ch 4].
where is the system’s Hamiltonian, including the environment’s perturbation. is the term responsible for decoherence to the environment.
@@
Although decoherence often occurs in femtoseconds or even sub-femtoseconds, specific molecular systems exhibit surprisingly long-lived coherence effects. For instance, Kaufman et al. [34] demonstrated that electronic coherences could survive for hundreds of femtoseconds in molecules with near parallel potential energy surfaces due to reduced dephasing between branched wavepackets (see also 4). Schultz et al. [16] comprehensively review cases where vibronic coupling and structured environments lead to long coherence lifetimes, including examples from light-harvesting complexes in photosynthesis.
Babcock et al. [35] reported superradiant effects in tryptophan networks extending over microtubule-scale architectures, which require long-lived electronic coherence. This is remarkable, since superradiance typically unfolds over hundreds of femtoseconds [36]. For isolated tryptophan in aqueous solution, electronic decoherence occurs in less than 100 fs [37], implying that the protein scaffold must be critical in preserving coherence to allow superradiant emission. Babcock’s result has also attracted media attention (see here and here) because it may support Hameroff-Penrose’s hypothesis on the quantum origin of consciousness (see Section 5.6.2).
@@
In addition to suppressing quantum interferences, decoherence also addresses the problem of the preferred basis. It leads to the selection of a stable outcome basis through a process known as einselection, which drives the quantum state to the basis least entangled with the environment (this basis is known as the pointer basis) [27]. This preferred basis depends on the Hamiltonian [39]. Assuming that the position operator commutes with the system-environment interaction Hamiltonian (usually the case), decoherence tends to select the position eigenstates as the preferred basis as long as the interaction Hamiltonian is more important for the dynamics than the system’s Hamiltonian. If both terms compete, the preferred basis is one close to coherent states, which minimizes the uncertainty in both position and momentum. If the system’s Hamiltonian dominates the dynamics, energy eigenstates are preferred as the basis.
For a molecule in a gas, Paz and Zurek’s analysis [39] implies that the outcome basis of the environment selects for measuring nuclei is approximately position states. On the other hand, energy states emerge as the preferred outcome basis for measuring electrons. Thus, when we initially chose the basis for describing the quantum state of the ammonia molecule, we did it because it matched our classical intuition of nuclei corresponding to particles with defined positions. After discussing decoherence, we realize that we may have developed such a classic intuition because the environment selects the basis and suppresses superposition effects, creating a world of particles with well-defined positions.
Nevertheless, einselection may not be the last word in the preferred basis problem. Adil and co-authors have recently shown that the preferred basis is not solely dictated by the system-environment interaction, but also by the natural factorization of the global Hilbert space [40]. Different factorizations can lead to multiple coexisting sets of pointer states, meaning classical behavior emerges in different ways depending on the structure of the Hamiltonian. These coexisting classical realms challenge the idea that einselection always yields a single classical outcome. They may even reignite the preferred basis problem, but now at the level of Hamiltonian factorization.
Decoherence explains why quantum superpositions become locally invisible; however, it does not explain why multiple observers consistently agree on what they observe. Zurek proposed that this consensus arises through Quantum Darwinism [41, 42], a process in which the environment acts not as a sink of information but as a communication channel. As the system interacts with its surroundings, information about its pointer states—those least perturbed by decoherence—is redundantly imprinted across many fragments of the environment. Different observers can then independently access these fragments to infer the same system properties without disturbing them. Objectivity, in this view, emerges as a Darwinian selection of information: only the most robust, redundantly recorded observables survive to be classically shared.
@@
This dominant role of decoherence in the current understanding of quantum mechanical processes does not imply a consensus. Landsmann, for instance, characterizes decoherence as “an unmitigated disaster” [43, Ch 11]. His core criticism is that decoherence rigorously only occurs in an environment with infinite degrees of freedom after infinite interaction time with the system. Kastner reinforces this criticism by noticing that decoherence claims of describing irreversible processes suffer from similar conceptual problems already present in statistical mechanics when explaining the emergence of irreversibility from reversible processes [44].
@@ 5 Facing the Problem of Outcomes
In the previous section, we saw that decoherence clarifies a significant portion of the Measurement Problem. However, the problem of outcomes is still open: After the basis is chosen and quantum superpositions are suppressed, the system remains in a mixture of possible outcomes. Decoherence does not tell how and why only one of these outcomes is measured. At this point, standard interpretation of quantum mechanics postulates an additional dynamic step, the collapse:
We illustrate these different processes with an example of a photoexcited molecule undergoing internal conversion in Figure 4. Light initially promotes the molecule to an excited electronic state. The nuclear wavepacket evolves on the potential energy surface of the excited state until it reaches a region of degeneracy with the ground state potential energy surface. There, quantum coherence builds between the two states, creating a state superposition. The dephasing between the wavepacket components in each state counteracts, causing decoherence. All events up to this point, including decoherence, can be simulated by propagating the Schrödinger equation for the molecule [45]. If the molecule is measured, it will be found to be either in the ground or excited state. However, our simulation can only tell the probabilities of these outputs.
@@
In the following subsections, we will dive into a few classes of potential solutions that represent the current debate on the Measurement Problem, with little attention to philosophical and historical aspects. More comprehensive approaches can be found in Refs.[4], [43, Introduction & Ch 11], [47, 48].
@@
As stated in Eq. (25), decoherence acts on the off-diagonal terms of the system’s reduced density matrix. Collapse, however, zeroes all diagonal terms except one in the entire density matrix—that is, the matrix describing the system, apparatus, and observer collectively, which Eq. (25) does not represent. We can be a bit more flexible with this definition. In principle, a quantum theory including collapse could go as far as to provide a diagonal density matrix. This means we would be left in a situation of statistical uncertainty: the collapse occurred, but the theory only informed us of the probability of each outcome. We could call such a case an epistemic collapse. If, however, the theory zeroes the diagonal (except one) elements, then we would have an ontic collapse.